Could you please solve this problem for me so I can explain it to my granddaughter?
3/b+7=13
Thank You

Answer by scott8148(6628)

subtracting 7 ___ 3/b = 6

multiplying by b ___ 3 = 6b

dividing by 6 ___ 3/6 = b ___ 1/2 = b